# Packing subgraphs in regular graphs

### v1

@@ 8 Concluding remarks
Recall that our main result (Theorem˜1.4) states that for every bipartite graph $H$, every dense regular graph $G$ contains an $H$-packing covering all but $O(1)$ vertices. As noted in the introduction, this was already known for unbalanced bipartite graphs $H$ by Kühn and Osthus [31] (see Theorem˜1.3). In fact, they proved a slightly stronger statement in this case: the host graph $G$ need only be almost regular rather than regular. More precisely, for every bipartite $H$ and every $c>0$, there exist $\varepsilon>0$ and $C$ such that any $n$-vertex graph $G$ in which every vertex has degree between $(c-\varepsilon)n$ and $(c+\varepsilon)n$ contains an $H$-packing covering all but at most $C$ vertices. This fails, however, for balanced bipartite $H$ (for example, if $G$ is a slightly unbalanced bipartite graph) so our result cannot, in general, be extended to the almost regular setting.
@@
It is plausible that, analogous to Theorem˜1.4, the conclusion holds for all tripartite $3$-graphs $H$ when $G$ is assumed to be regular.
@@
Is it true that for every tripartite $3$-graph $H$ and every $0<c<1$, there exist $\varepsilon>0$ and $C>0$ such that the following holds? If $G$ is a $3$-graph on $n$ vertices in which all vertices have the same degree at least $cn^{2}$, and every pair of vertices has degree at least $cn$, then $G$ contains an $H$-packing covering all but at most $C$ vertices.
The problem of partitioning the vertex set of edge-coloured complete graphs into a small number of monochromatic subgraphs has a very rich history; see [12] for a recent survey. An early example of a problem of this kind is Lehel’s conjecture. An interesting problem in this direction, suggested by Matija Bucić, is to consider a variant of Theorem˜1.4 where the edges of $G$ are colored and we seek packings using monochromatic $H$-copies (where different copies of $H$ may receive different colors).
@@
Is it true that for every bipartite graph $H$, integer $r\geq 2$, and $0<c<1$, there exists $C>0$ such that the following holds? If $G$ is an $n$-vertex $d$-regular graph with $d\geq cn$, whose edges are colored with $r$ colors, then $G$ contains a collection of vertex-disjoint monochromatic copies of $H$ covering all but at most $C$ vertices.
@@ Appendix B Hamiltonicity of clusters
In this section, we present the proof of Lemma˜3.3. Our proof closely follows that of Lemma 4 in [11], but we include it here for completeness and to address differences in the lemma’s statement.
@@
We shall use the following corollary of Theorem˜B.1.
@@
Given vertices $x$ and $y$, form $G^{\prime}$ by adding the edge $xy$ to $G$, removing the edge $yx$ (if it exists), and removing all the edges directed toward $y$ or away from $x$. Next, form $G^{\prime\prime}$ by contracting the edge $xy$. It is easy to check that $G^{\prime\prime}$ is a robust $(\nu/2,2\tau)$-out-expander. Thus, by Theorem˜B.1, it contains a Hamilton cycle. This cycle corresponds to a Hamilton cycle in $G^{\prime}$ which contains the edge $xy$, which in turn corresponds to a Hamilton path in $G$ with ends $x$ and $y$. ∎
@@
The following claim will allow us to use Corollary˜B.2 above.
@@
It follows from Claim˜B.3 that $H$ is a robust $(\xi,2\xi^{1/7})$-expander. Let $x,y\in V$, where $x\in X$ and $y\in Y$ if $G$ is bipartite. Our goal is to show that $H$ contains a Hamilton path with ends $x$ and $y$. First, consider the case where $G$ is $\gamma$-far-from-bipartite. Form a digraph $D$ by replacing each edge $uv$ of $H$ with the two directed edges $uv$ and $vu$. It follows from Claim˜B.3 that $D$ is a robust $(\xi,2\xi^{1/7})$-out-expander. Then, Corollary˜B.2 implies the existence of a Hamilton path in $D$ with ends $x$ and $y$, which corresponds to a Hamilton path in $H$ with the same ends.
Now, suppose that $G$ is bipartite. Note that in this case $H$ is a balanced bipartite graph (by our choice of the set $W$). We first show that $H$ has a perfect matching by verifying Hall’s condition i.e., we will show that for every $S\subseteq X^{\prime}$, we have $|N_{H}(S)|\geq|S|$. Notice that $H$ has minimum degree at least $\zeta(n-1)-\xi n\geq\zeta n/2$, by the assumption that $G$ has no $\zeta$-sparse cuts and by $|W|\leq\xi n$. Thus, if $|S|<\zeta n/2$ (and $S\neq\emptyset$) then $|N_{H}(S)|\geq\zeta n/2\geq|S|$. Similarly, if $|S|>|V|/2-\zeta n/2$ then $|N_{H}(S)|=|V|/2\geq|S|$. Hence, we may assume that $\zeta n/2\leq|S|\leq|V|/2-\zeta n/2$, which implies that $\xi^{1/7}n\leq|S|\leq(1/2-\xi^{1/7})n$, but then Claim˜B.3 implies that $|N_{H}(S)|\geq|S|$, showing that $H$ has a perfect matching.
Let $\{a_{1}b_{1},\ldots,a_{t}b_{t}\}$ be a perfect matching in $H$, where $t=|X^{\prime}|$ and $a_{i}\in X^{\prime},b_{i}\in Y^{\prime}$ for $i\in[t]$. We assume for convenience that $a_{i}b_{i}$ is not the edge $xy$ (if the edge $xy$ exists) for $i\in[t]$ – this is possible because removal of the edge $xy$ from $H$ does not affect the arguments above. Without loss of generality, $a_{1}=x$ and $b_{t}=y$. Form an auxiliary directed graph $D$ with vertex set $\{v_{1},\ldots,v_{t}\}$ where $v_{i}v_{j}$ is a directed edge whenever $b_{i}a_{j}$ is an edge of $H$. It follows from Claim˜B.3 that $D$ is a robust $(\xi,3\xi^{1/7})$-out-expander. Then, by Corollary˜B.2, there is a Hamilton path in $D$ starting at $v_{1}$ and ending at $v_{t}$. Without loss of generality, suppose that this path is $v_{1},\ldots,v_{t}$. This path corresponds to the Hamilton path $x=a_{1},b_{1},\ldots,a_{t},b_{t}=y$ in $H$, as desired. ∎

### v2

@@ 8 Concluding remarks
Recall that our main result (Theorem 1.4) states that for every bipartite graph $H$, every dense regular graph $G$ contains an $H$-packing covering all but $O(1)$ vertices. As noted in the introduction, this was already known for unbalanced bipartite graphs $H$ by Kühn and Osthus [31] (see Theorem 1.3). In fact, they proved a slightly stronger statement in this case: the host graph $G$ need only be almost regular rather than regular. More precisely, for every bipartite $H$ and every $c>0$, there exist $\varepsilon>0$ and $C$ such that any $n$-vertex graph $G$ in which every vertex has degree between $(c-\varepsilon)n$ and $(c+\varepsilon)n$ contains an $H$-packing covering all but at most $C$ vertices. This fails, however, for balanced bipartite $H$ (for example, if $G$ is a slightly unbalanced bipartite graph) so our result cannot, in general, be extended to the almost regular setting.
@@
It is plausible that, analogous to Theorem 1.4, the conclusion holds for all tripartite $3$-graphs $H$ when $G$ is assumed to be regular.
@@
Is it true that for every tripartite $3$-graph $H$ and every $0<c<1$, there exist $\varepsilon>0$ and $C>0$ such that the following holds? If $G$ is a $3$-graph on $n$ vertices in which all vertices have the same degree, which is at least $cn^{2}$, and every pair of vertices has degree at least $cn$, then $G$ contains an $H$-packing covering all but at most $C$ vertices.
The problem of partitioning the vertex set of edge-coloured complete graphs into a small number of monochromatic subgraphs has a very rich history; see [12] for a recent survey. An early example of a problem of this kind is Lehel’s conjecture. An interesting problem in this direction, suggested by Matija Bucić, is to consider a variant of Theorem 1.4 where the edges of $G$ are coloured and we seek packings using monochromatic $H$-copies (where different copies of $H$ may receive different colours).
@@
Is it true that for every bipartite graph $H$, integer $r\geq 2$, and $0<c<1$, there exists $C>0$ such that the following holds? If $G$ is an $n$-vertex $d$-regular graph with $d\geq cn$, whose edges are coloured with $r$ colours, then $G$ contains a collection of vertex-disjoint monochromatic copies of $H$ covering all but at most $C$ vertices.
@@ Appendix B Hamiltonicity of clusters
In this section, we present the proof of Lemma 3.3. Our proof closely follows that of Lemma 4 in [11], but we include it here for completeness and to address differences in the lemma’s statement.
@@
We shall use the following corollary of Theorem B.1.
@@
Given vertices $x$ and $y$, form $G^{\prime}$ by adding the edge $xy$ to $G$, removing the edge $yx$ (if it exists), and removing all the edges directed toward $y$ or away from $x$. Next, form $G^{\prime\prime}$ by contracting the edge $xy$. It is easy to check that $G^{\prime\prime}$ is a robust $(\nu/2,2\tau)$-out-expander. Thus, by Theorem B.1, it contains a Hamilton cycle. This cycle corresponds to a Hamilton cycle in $G^{\prime}$ which contains the edge $xy$, which in turn corresponds to a Hamilton path in $G$ with ends $x$ and $y$. ∎
@@
The following claim will allow us to use Corollary B.2 above.
@@
It follows from Claim B.3 that $H$ is a robust $(\xi,2\xi^{1/7})$-expander. Let $x,y\in V$, where $x\in X$ and $y\in Y$ if $G$ is bipartite. Our goal is to show that $H$ contains a Hamilton path with ends $x$ and $y$. First, consider the case where $G$ is $\gamma$-far-from-bipartite. Form a digraph $D$ by replacing each edge $uv$ of $H$ with the two directed edges $uv$ and $vu$. It follows from Claim B.3 that $D$ is a robust $(\xi,2\xi^{1/7})$-out-expander. Then, Corollary B.2 implies the existence of a Hamilton path in $D$ with ends $x$ and $y$, which corresponds to a Hamilton path in $H$ with the same ends.
Now, suppose that $G$ is bipartite. Note that in this case $H$ is a balanced bipartite graph (by our choice of the set $W$). We first show that $H$ has a perfect matching by verifying Hall’s condition i.e., we will show that for every $S\subseteq X^{\prime}$, we have $|N_{H}(S)|\geq|S|$. Notice that $H$ has minimum degree at least $\zeta(n-1)-\xi n\geq\zeta n/2$, by the assumption that $G$ has no $\zeta$-sparse cuts and by $|W|\leq\xi n$. Thus, if $|S|<\zeta n/2$ (and $S\neq\emptyset$) then $|N_{H}(S)|\geq\zeta n/2\geq|S|$. Similarly, if $|S|>|V|/2-\zeta n/2$ then $|N_{H}(S)|=|V|/2\geq|S|$. Hence, we may assume that $\zeta n/2\leq|S|\leq|V|/2-\zeta n/2$, which implies that $\xi^{1/7}n\leq|S|\leq(1/2-\xi^{1/7})n$, but then Claim B.3 implies that $|N_{H}(S)|\geq|S|$, showing that $H$ has a perfect matching.
Let $\{a_{1}b_{1},\ldots,a_{t}b_{t}\}$ be a perfect matching in $H$, where $t=|X^{\prime}|$ and $a_{i}\in X^{\prime},b_{i}\in Y^{\prime}$ for $i\in[t]$. We assume for convenience that $a_{i}b_{i}$ is not the edge $xy$ (if the edge $xy$ exists) for $i\in[t]$ – this is possible because removal of the edge $xy$ from $H$ does not affect the arguments above. Without loss of generality, $a_{1}=x$ and $b_{t}=y$. Form an auxiliary directed graph $D$ with vertex set $\{v_{1},\ldots,v_{t}\}$ where $v_{i}v_{j}$ is a directed edge whenever $b_{i}a_{j}$ is an edge of $H$. It follows from Claim B.3 that $D$ is a robust $(\xi,3\xi^{1/7})$-out-expander. Then, by Corollary B.2, there is a Hamilton path in $D$ starting at $v_{1}$ and ending at $v_{t}$. Without loss of generality, suppose that this path is $v_{1},\ldots,v_{t}$. This path corresponds to the Hamilton path $x=a_{1},b_{1},\ldots,a_{t},b_{t}=y$ in $H$, as desired. ∎